# DPO Meets PPO: Reinforced Token Optimization for RLHF

### v2

@@ Abstract
In the classical Reinforcement Learning from Human Feedback (RLHF) framework, Proximal Policy Optimization (PPO) is employed to learn from sparse, sentence-level rewards—a challenging scenario in traditional deep reinforcement learning. Despite the great successes of PPO in the alignment of state-of-the-art closed-source large language models (LLMs), its open-source implementation is still largely sub-optimal, as widely reported by numerous research studies. To address these issues, we introduce a framework that models RLHF problems as a Markov decision process (MDP), enabling the capture of fine-grained token-wise information. Furthermore, we provide theoretical insights that demonstrate the superiority of our MDP framework over the previous sentence-level bandit formulation. Under this framework, we introduce an algorithm, dubbed as Reinforced Token Optimization (RTO), which learns the token-wise reward function from preference data and performs policy optimization based on this learned token-wise reward signal. Theoretically, RTO is proven to have the capability of finding the near-optimal policy sample-efficiently. For its practical implementation, RTO innovatively integrates Direct Preference Optimization (DPO) and PPO. DPO, originally derived from sparse sentence rewards, surprisingly provides us with a token-wise characterization of response quality, which is seamlessly incorporated into our subsequent PPO training stage. Extensive real-world alignment experiments verify the effectiveness of the proposed approach.
@@ 1 Introduction
Despite the success of this framework in the aforementioned powerful closed-source LLMs, the training of PPO is known to be unstable and sample-inefficient (Choshen et al., 2019) compared to supervised learning. PPO frequently fails to maintain a consistent average response length or experiences sudden drops in reward value. Moreover, the superior performance of PPO also relies on the code-level optimization and an appropriate configuration of the hyper-parameters (Engstrom et al., 2020), while the training stability issue further prohibits us from achieving the best performance of the PPO. So far, the success of PPO has not been widely reproduced, especially in the open-source community with rather limited resources. While researchers have made efforts to propose alternative approaches to the PPO algorithm, with notable examples like rejection sampling fine-tuning (Dong et al., 2023; Gulcehre et al., 2023), direct preference learning algorithms (Rafailov et al., 2023; Zhao et al., 2023; Azar et al., 2023), there is little evidence that these newly proposed approaches alone can make the state-of-the-art LLMs. Therefore, improving the performance of the PPO algorithm in the context of RLHF is still an important research direction that is largely under-explored.
After examining the open-source implementation of PPO, we identify that one potential reason for the sub-optimal performance of PPO is the mismatch between the formulation of RLHF and the nature of PPO. Specifically, in the existing framework (Ouyang et al., 2022; Bai et al., 2022), RLHF is formulated as a bandit, where the entire response sentence is considered to be an action, and the reward is sentence-level, evaluating only the overall quality of the response. However, PPO is designed for multi-step RL problems modeled as Markov decision processes (MDPs), requiring a token-wise reward assignment to each step. In typical implementations of PPO (e.g., the TRL package from huggingface [^1] https://github.com/huggingface/trl), besides the regularization reward function assigned to each token to ensure the fine-tuned LLM stays close to the supervised fine-tuning (SFT) model, the learned sentence-level reward is only distributed to the last token, while other tokens receive zero learned reward. See (2.3) for the formal mathematical description. Clearly, there is a separation in terms of the assignment strategies of the regularization reward and the learned reward. Meanwhile, while it is generally believed that a fine-grained characterization with token-wise feedback can provide more information, in practice, it is also challenging to collect effective token-wise feedback for human conversations and use it in the MLE process. Consequently, the construction of token-wise reward signals also remains largely under-explored in the literature of RLHF.
@@ 1.1 Our Contributions
  • Moving toward the practical implementation of RTO, we adopt a novel token-wise reward extraction approach from direct preference optimization (DPO; Rafailov et al., 2023). By assigning this DPO-based token-wise reward function to each token and then optimizing with PPO, RTO outperforms existing baselines such as PPO and DPO in the task of dialogue.
@@ 1.2 Related Works
Although some works (e.g., Uesato et al., 2022; Lightman et al., 2023; Yang et al., 2024b) use token-wise or step-wise information to enhance the performance of LLMs, such as their reasoning ability, we will not discuss them in detail here. Instead, we will focus on comparing our work with others that aim to improve the PPO in RLHF. In particular, Li et al. (2023a) and Ahmadian et al. (2024) state that the PPO is not the best fit for RLHF because of the sentence-level reward and deterministic transition, and argue that the reinforce-style (Williams, 1992) algorithms perform better. Wu et al. (2024) proposes to construct several separate reward functions for different goals and use the linear combination of them to guide the PPO training, but the separate models are still confined to the sentence level. Similarly, Jang et al. (2023) extends the PPO to the multi-objective optimization scenario, but still uses the sentence-level modeling. Chan et al. (2024) shares similar insights that aim to improve PPO via a dense reward. They still follow the two-staged RLHF framework to model the reward function via MLE of the Bradley-Terry model and assume that the learned reward is based on the transformer (Vaswani et al., 2017). Then, they propose to use the attention value to redistribute the final scalar reward on a token level. In comparison, while sharing similar insights about using a token-wise reward, our techniques to obtain the dense signal and mathematical motivation are fundamentally different.
Concurrent work.
During the preparation of this work, there is a concurrent and independent work (Rafailov et al., 2024) that also provides a token-wise MDP formulation for RLHF. Their work shares the same insight as ours, namely that “DPO implicitly optimizes the token-wise reward”. Based on this insight, they improve the efficiency of search-based algorithms. In contrast, we propose a new algorithm RTO that leverages the token-wise reward functions to enhance the performance of PPO. In addition, our work provides a theoretical foundation for the unique advantages of token-wise MDP and its sample-efficient learning.
@@ 4.1 Theoretical Version with Sample Complexity Guarantee
There have also been previous works (Pacchiano et al., 2021; Chen et al., 2022; Wang et al., 2023; Li et al., 2023b; Zhan et al., 2023a) studying RLHF under the MDP framework, also known as dueling RL and preference-based RL. However, these works do not consider the KL constraint, which is an essential component of RLHF. Furthermore, they do not explicitly emphasize the superiority of the MDP framework over the contextual dueling bandit problem in the context of LLMs, and their proposed algorithms lack practical implementation. In contrast, we will provide a practical implementation of our algorithm, demonstrating the practicality of our approach.
@@
6:     Calculate the token-wise reward $r_{\mathrm{rto}}$ for each pair $(x,y)$ by $\triangleright$ token-wise reward learning
$r_{\mathrm{rto}}((x,y_{1:h-1}),y_{h})=\beta_{1}\log\frac{\pi_{\mathrm{dpo}}(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}-\beta_{2}\log\frac{\pi_{t-1}(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}.$
@@ 4.2 Practical Implementation
$\beta_{1}\log\frac{\pi_{\beta}^{*}(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}-\beta_{2}\log\frac{\pi(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}$ (4.6)
$\qquad\approx\beta_{1}\log\frac{\pi_{\mathrm{dpo}}(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}-\beta_{2}\log\frac{\pi(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}:=r_{\mathrm{rto}}((x,y_{1:h-1}),y_{h})$
as the token-wise reward used by RTO, where $\beta_{1}$ and $\beta_{2}$ are tuning parameters, and $\pi$ is the current policy to be updated. In the last step of (LABEL:eq:prac:5), we use $\pi_{\mathrm{dpo}}$, the policy learned by DPO, as a proxy for the unknown $\pi_{\beta}^{*}$. Finally, we employ PPO to optimize the token-wise reward $r_{\mathrm{rto}}$ in (LABEL:eq:prac:5).
@@ 5 Experiments
In this section, we conduct real-world alignment experiments to verify the effectiveness of RTO. We provide the experimental setups and experimental results in Sections 5.1 and 5.2, respectively.
5.1 Experimental Setups
Tasks and Data.
We consider two text generation tasks, including the Dialogue Generation (Bai et al., 2022) and Summarization task (Völske et al., 2017). Given a text sequence ($x$) representing dialogue history between the user and the assistant, the goal of the Dialogue Generation task is to generate a helpful response ($y$) as the answer. For this purpose, we utilize the helpful subset of the Anthropic Helpful and Harmless (HH-RLHF) dialogue dataset [^3] https://huggingface.co/datasets/Anthropic/hh-rlhf (Bai et al., 2022). Each sample of the HH-RLHF dataset is accompanied by a history and two alternative responses, with preferences annotated by humans. For the Summarization task, the model is required to generate a concise summary for a given post from the Reddit forum. Specifically, we self-supervise fine-tune the foundational model using the Reddit TL;DR summarization dataset (Völske et al., 2017), where each data point comprises a post $x$ and its corresponding summary $y$. Subsequently, we align the model with human preferences using its preference version [^4] https://huggingface.co/datasets/openai/summarize_from_feedback, where each data point comprises a post and two summaries, with preferences annotated by humans. To facilitate readers, we provide examples of the HH-RLHF and TL;DR datasets in Appendix D.1.
Model and Baselines.
We employ the open-sourced Pythia-2.8B model (Biderman et al., 2023) as the backbone for all experiments. We select four methods as baselines. The first baseline, SFT (i.e., $\pi_{\mathrm{ref}}$), fine-tunes the language model using the human-preferred responses only. Building upon this SFT model, we further train a DPO model, which finetunes the SFT model using the positive/negative preference data. Besides these two RL-free algorithms, we compare two RLHF algorithms relying on RL training. The first one is the standard PPO algorithm, which directly optimizes sentence-level reward in (2.3) from the SFT model. For an ablation study, we also consider a PPO variant as a baseline where the sentence-level reward is provided by the DPO objective, defined as follows:
${r}_{\mathrm{dppo}}(x,y_{1:h})=\begin{cases}{{0}}-\beta_{2}\log\frac{\pi(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}&\text{ if }h\leq H-1,\\ {\beta_{1}\log\frac{\pi(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}}-\beta_{2}\log\frac{\pi(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}&\text{ if }h=H,\end{cases}$ (5.1)
where $(x,y=y_{1:H})$ is the prompt-response pair, $\pi$ is the current policy, and $(\beta_{1},\beta_{2})$ are tuning hyperparameters. In other words, we use the DPO to extract a sentence-level reward from the preference data, assign it to the last token, and fine-tune the model using the PPO algorithm. We refer to this baseline as DPPO. For our proposed RTO, we use the DPO model to derive a token-wise reward model (LABEL:eq:prac:5), and train the policy to align human preference using PPO, as detailed in Algorithm 2. The training configurations of all aforementioned models are given in Appendix D.2.
Evaluation.
We primarily assess the alignment performance of various methods using GPT-4. The GPT-4 evaluation harnesses the capabilities of GPT-4 itself and has been shown to align well with human evaluations (Rafailov et al., 2023). For the same prompt, we provide GPT-4 with two responses generated by two different models and ask it to determine which one is superior. We then calculate the win rates, following Rafailov et al. (2023). The prompts for GPT-4 evaluation of both HH-RLHF and TL;DR are presented in Appendix D.3. For each GPT-4 evaluation, we use 100 samples. In addition to GPT-4 evaluation, we also employ oracle reward evaluation for HH-RLHF. For this, we employ an open-sourced reward model  [^5] https://huggingface.co/weqweasdas/RM-Mistral-7B as the oracle, which is trained from Mistral-7B (Jiang et al., 2023) and achieves one of the highest accuracy on the Anthropic Helpful and Harmless dialogue task (Bai et al., 2022). Given 400 prompts, we generate the responses using all baselines and calculate the average rewards of the responses by the oracle reward model.
5.2 Experimental Results
Win Rates.
The experiment results of our proposed method and baselines on HH-RLHF are detailed in Table 1(a) and Figure 3(c). This table meticulously presents the win rates between different models, assessed through both the oracle model and the GPT-4.
Win Rate RTO DPO SFT PPO DPPO
RTO 0.50 0.51 0.56 0.52 0.56
DPO 0.49 0.50 0.52 0.54 0.51
SFT 0.44 0.48 0.50 0.41 0.46
PPO 0.48 0.46 0.59 0.50 0.52
DPPO 0.44 0.49 0.54 0.48 0.50
Table 1(a): HH-RLHF win rates evaluated by GPT-4.
Win Rate RTO DPO SFT PPO DPPO
RTO 0.50 0.61 0.67 0.67 0.67
DPO 0.39 0.50 0.58 0.59 0.50
SFT 0.33 0.42 0.50 0.59 0.49
PPO 0.33 0.41 0.41 0.50 0.40
DPPO 0.33 0.50 0.51 0.60 0.50
Table 1(b): TL;DR win rates evaluated by GPT-4
((a)) HH-RLHF Evaluated by GPT-4
((b)) TL;DR Evaluated by GPT-4
((c)) HH-RLHF Evaluated by oracle model
Figure 3: Subfigures (a) and (b) show the win rates of RTO across different sampling temperatures, while Subfigure (c) visualizes the mean reward of different baselines evaluated by the oracle model.
From these results, we can see that the model trained by RTO achieves win rates over $50\%$ against all other baselines, especially compared to DPO, evaluated by both the oracle reward model and GPT-4. This highlights the effectiveness of the RTO algorithm in alignment tasks. Furthermore, the model trained by RTO gets a win rate of $61.1\%$ evaluated by the oracle reward model and a win rate of $56\%$ evaluated by GPT-4 over the DPPO algorithm. This implies that the token-wise reward mechanism significantly improves the performance of the RL algorithm in training models. Table 1(b) presents the performance of our method on the TL;DR dataset. We can see that the model trained by RTO outperforms all other baselines. Specifically, we achieve a win rate of $61\%$ over the DPO algorithm evaluated by GPT-4. This illustrates the effectiveness of the RTO algorithm in a real-world text summarization task. All these empirical findings demonstrate the token-wise reward mechanism’s advantage in improving model performance.
Ablation Studies on Temperatures.
We further compare the model trained by RTO to other baselines on both datasets across different temperatures. In Figure 3(c), we present the results for these methods as we vary the temperature. We can observe that the RTO model consistently demonstrates superior performance compared to other baselines, highlighting its robustness across different temperatures.
Optimization Process Curves.
To further investigate the benefits of the token-wise reward mechanism in the optimization process, we compare the estimated reward during the training period in Figure 4 and Figure 4. In both figures, the x-axis represents the training iterations. The y-axis represents the reward given by the implicit reward model derived from the DPO model (the reward model used in training) per batch. As we can see, in one epoch (roughly corresponds to 160 PPO training iterations in Figure 4 and 240 in Figure 4), the reward of the model trained by RTO on HH-RLHF can achieve about $0.1$, while the reward of the model trained by DPPO is roughly $-0.25$; similarly, the reward of the model trained by RTO on TL;DR can achieve about $0.4$, while the reward of the model trained by DPPO is roughly $-0.2$. The results demonstrate that the token-wise reward mechanism significantly enhances the training process, leading to a remarkably higher reward.
Figure 4: The reward curve of DPPO and RTO during training. The reward is given by the implicit reward model $\beta\log\frac{\pi_{\mathrm{dpo}}(y{\,|\,}x)}{\pi_{\mathrm{ref}}(y{\,|\,}x)}$ optimized by DPO. The x-axis represents the training steps, and the y-axis represents the reward values. Figure  4 shows the reward curve of HH-RLHF, and Figure  4 shows the reward curve of TL;DR.

### v3

@@ Abstract
In the classical Reinforcement Learning from Human Feedback (RLHF) framework, Proximal Policy Optimization (PPO) is employed to learn from sparse, sentence-level rewards—a challenging scenario in traditional deep reinforcement learning. Despite the great successes of PPO in the alignment of large language models, its open-source implementation is still largely sub-optimal. To address these issues, we introduce a framework that models RLHF problems as a Markov decision process (MDP), enabling the capture of fine-grained token-wise information. Under this framework, we introduce an algorithm Reinforced Token Optimization (RTO), which learns the token-wise reward function from preference data and performs policy optimization based on this learned token-wise reward signal.Theoretically, RTO is proven to have the capability of finding the near-optimal policy sample-efficiently. For its practical implementation, RTO innovatively integrates Direct Preference Optimization (DPO) and PPO. DPO, originally derived from sparse sentence rewards, surprisingly provides us with a token-wise characterization of response quality, which is seamlessly incorporated into our subsequent PPO training stage. Extensive experiments demonstrate that RTO performs better than PPO and other direct preference learning algorithms. In particular, RTO outperforms PPO by 7.5 points on the AlpacaEval 2 benchmark and by 4.1 points on Arena-Hard. Our code and models are available at https://github.com/zkshan2002/RTO.
@@ 1 Introduction
Despite the success of this framework in the aforementioned powerful closed-source LLMs, the training of PPO is known to be unstable and sample-inefficient (Choshen et al., 2019). While researchers have made efforts to propose alternative approaches to the PPO algorithm, with notable examples like rejection sampling fine-tuning (Dong et al., 2023; Gulcehre et al., 2023), direct preference learning algorithms (Rafailov et al., 2023; Zhao et al., 2023; Azar et al., 2023), there is little evidence that these newly proposed approaches alone can make the state-of-the-art LLMs. Therefore, improving the performance of the PPO algorithm in the context of RLHF is still an important research direction that is largely under-explored.
After examining the open-source implementation of PPO, we identify that one potential reason for the sub-optimal performance of PPO is the mismatch between the formulation of RLHF and the nature of PPO. Specifically, in the existing framework (Ouyang et al., 2022; Bai et al., 2022), RLHF is formulated as a bandit, where the entire response sentence is considered to be an action, and the reward is sentence-level, evaluating only the overall quality of the response. However, PPO is designed for multi-step RL problems modeled as Markov decision processes (MDPs), requiring a token-wise reward assignment to each step. In typical implementations of PPO (e.g., OpenRLHF and TRL), besides the regularization reward function assigned to each token to ensure the fine-tuned LLM stays close to the supervised fine-tuning (SFT) model, the learned sentence-level reward is only distributed to the last token, while other tokens receive zero learned reward. See (2.3) for the formal mathematical description. Clearly, there is a separation in terms of the assignment strategies of the regularization reward and the learned reward. Meanwhile, while it is generally believed that a fine-grained characterization with token-wise feedback can provide more information, in practice, it is also challenging to collect effective token-wise feedback for human conversations and use it in the MLE process. Consequently, the construction of token-wise reward signals also remains largely under-explored in the literature of RLHF.
@@ 1.1 Our Contributions
  • Moving toward the practical implementation of RTO, we adopt a novel token-wise reward extraction approach from direct preference optimization (DPO; Rafailov et al., 2023). By assigning this DPO-based token-wise reward function to each token and then optimizing with PPO. RTO demonstrates superior performance compared to both PPO and direct preference learning baselines such as DPO (Rafailov et al., 2023), R-DPO (Park et al., 2024), and SimPO (Meng et al., 2024). In particular, RTO achieves a 7.5-point improvement on the AlpacaEval 2 benchmark and a 4.1-point improvement on Arena-Hard. Additionally, RTO exhibits strong data scaling properties compared to PPO — (i) reaching PPO-level performance with only $1/8$ of the data and (ii) continuing to improve as more data is added, whereas PPO saturates early.
@@ 1.2 Related Works
Although some works (e.g., Uesato et al., 2022; Lightman et al., 2023; Yang et al., 2024b) use token-wise or step-wise information to enhance the performance of LLMs, such as their reasoning ability, we will not discuss them in detail here. Instead, we will focus on comparing our work with others that aim to improve the PPO in RLHF. In particular, Li et al. (2023b) and Ahmadian et al. (2024) state that the PPO is not the best fit for RLHF because of the sentence-level reward and deterministic transition, and argue that the reinforce-style (Williams, 1992) algorithms perform better. Wu et al. (2024) proposes to construct several separate reward functions for different goals and use the linear combination of them to guide the PPO training, but the separate models are still confined to the sentence level. Similarly, Jang et al. (2023) extends the PPO to the multi-objective optimization scenario, but still uses the sentence-level modeling. Chan et al. (2024) shares similar insights that aim to improve PPO via a dense reward. They still follow the two-staged RLHF framework to model the reward function via MLE of the Bradley-Terry model and assume that the learned reward is based on the transformer (Vaswani et al., 2017). Then, they propose to use the attention value to redistribute the final scalar reward on a token level. In comparison, while sharing similar insights about using a token-wise reward, our techniques to obtain the dense signal and mathematical motivation are fundamentally different.
Concurrent and Subsequent work.
We notice several concurrent and independent works by Rafailov et al. (2023); Zeng et al. (2024); Meng et al. (2024). Rafailov et al. (2024) also provides a token-wise MDP formulation for RLHF. Their work shares the same insight as ours, namely that “DPO implicitly optimizes the token-wise reward”. Based on this insight, they improve the efficiency of search-based algorithms. In contrast, we propose a new algorithm RTO that leverages the token-wise reward functions to enhance the performance of PPO. In addition, our work provides a theoretical foundation for the unique advantages of token-wise MDP and its sample-efficient learning. Meanwhile, Zeng et al. (2024) and Meng et al. (2024) introduce two direct preference learning algorithms (token-wise DPO and SimPO). Unlike these approaches, our focus is on improving PPO-based RL training by leveraging token-wise rewards. We include these two algorithms as baselines to demonstrate the superior performance of RTO. Finally, following our work, Cui et al. (2025); Yin et al. (2025) utilizes implicit rewards in RL training to enhance chat and reasoning capabilities, highlighting the broad applicability of our method.
@@ 4.1 Theoretical Version with Sample Complexity Guarantee
There have also been previous works (Pacchiano et al., 2021; Chen et al., 2022; Wang et al., 2023; Li et al., 2023c; Zhan et al., 2023a) studying RLHF under the MDP framework, also known as dueling RL and preference-based RL. However, these works do not consider the KL constraint, which is an essential component of RLHF. Furthermore, they do not explicitly emphasize the superiority of the MDP framework over the contextual dueling bandit problem in the context of LLMs, and their proposed algorithms lack practical implementation. In contrast, we will provide a practical implementation of our algorithm, demonstrating the practicality of our approach.
@@
6:     Calculate the token-wise reward $r_{\mathrm{rto}}$ for each pair $(x,y)$ by (4.7). $\triangleright$ token-wise reward learning
@@ 4.2 Practical Implementation
$\beta_{1}\log\frac{\pi_{\beta}^{*}(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}-\beta_{2}\log\frac{\pi(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}\approx\beta_{1}\log\frac{\pi_{\mathrm{dpo}}(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}-\beta_{2}\log\frac{\pi(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}$ (4.6)
as the token-wise reward, where $\beta_{1}$ and $\beta_{2}$ are tuning parameters, and $\pi$ is the current policy to be updated. In the last step of (4.6), we use $\pi_{\mathrm{dpo}}$, the policy learned by DPO, as a proxy for the unknown $\pi_{\beta}^{*}$. Finally, we employ PPO to optimize the following token-wise reward $r_{\mathrm{rto}}$
${r}_{\mathrm{rto}}(x,y_{1:h})=\begin{cases}\beta_{1}\log\frac{\pi_{\mathrm{dpo}}(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}-\beta_{2}\log\frac{\pi(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}&\text{ if }h\leq H-1,\\ \beta_{1}\log\frac{\pi_{\mathrm{dpo}}(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}-\beta_{2}\log\frac{\pi(y_{h}{\,|\,}x,y_{1:h-1})}{\pi_{\mathrm{ref}}(y_{h}{\,|\,}x,y_{1:h-1})}+\beta_{3}\cdot r_{\mathrm{MLE}}(x,y_{1:H})&\text{ if }h=H,\end{cases}$ (4.7)
where $\beta_{3}\geq 0$ is a tuning parameter and $r_{\mathrm{MLE}}$ represents a sentence-level reward. This additional sentence-level reward helps prevent responses from becoming either extremely long or extremely short. This aligns with the observation that ensemble rewards can effectively mitigate the overoptimization issues (Coste et al., 2023). We also remark that the sentence-level reward $r_{\text{MLE}}$ can be much smaller in magnitude compared to both the policy model (actor) and DPO reward model, making the overall computational cost of RTO comparable to the standard RLHF pipeline: The lower cost of using a much smaller critic in PPO compensates for both the small extra cost required by DPO than reward model, and the training and serving of the tiny reward model. For parameter selection, $\beta_{3}$ can be set to $1$ when $r_{\text{MLE}}$ is included. This choice is without loss of generality, as the key factor is the ratio of $\beta_{3}$ to $\beta_{1}$ and $\beta_{2}$, rather than its absolute value. $\beta_{2}$ can be chosen similarly in standard PPO configurations. The only extra hyperparameter $\beta_{1}$ can be set small to prevent the DPO reward from dominating, thereby requiring minimum tuning.
@@ 5 Experiments
In this section, we conduct comprehensive alignment experiments to verify the effectiveness of RTO.
5.1 Benchmark Results
We present a thorough comparison of RTO with PPO and other widely used direct preference learning algorithms on popular benchmarks to highlight RTO’s strong performance.
Task, Data, and Evaluation.
To assess the overall quality of generated text responses across multiple dimensions (e.g., helpfulness, accuracy, and clarity), we employ the dataset UltraFeedback (Cui et al., 2023) that contains comprehensive human feedback annotations on model outputs. We evaluate models using two established benchmarks: AlpacaEval 2 (Li et al., 2023a) and Arena-Hard (Li et al., 2024). These benchmarks assess various conversational abilities across different types of queries. For AlpacaEval 2, we report both standard win rates (WR) and length-controlled win rates (LC). For Arena-Hard, we present the WR along with its style-controlled (SC) version. Both LC and SC are specifically designed to mitigate verbosity bias of llm judge.
Implementation Details of RTO and Baselines.
We employ Llama-3-8B (Dubey et al., 2024) as the base model. For our comparative analysis, we implement several baselines. All subsequent models are initialized with an open-source SFT model Dong et al. (2024) that fine-tunes Llama-3-8B with a diverse mixture of high-quality data. We further train a DPO model, which finetunes the SFT model using the positive/negative preference data. Besides these two RL-free algorithms, we compare three RLHF algorithms relying on RL training. The first one is the standard PPO algorithm, which directly optimizes sentence-level 8B reward in (2.3). Our proposed RTO algorithm leverages both token-wise signals from the DPO model and an additional 1B sentence-wise reward $r_{\mathrm{MLE}}$ to compute the RTO reward specified in (4.7). The policy is then trained to align with human preferences using PPO updates, as detailed in Algorithm 2. For a comprehensive comparison, we include the length-controlled version of DPO, referred to as R-DPO (Park et al., 2024), along with two preference learning algorithms introduced in concurrent and independent works: SimPO (Meng et al., 2024) and token-wise DPO (TDPO) (Zeng et al., 2024), as baselines. We include more details in appendix D.
RTO Outperforms PPO and Other Direct Preference Learning Algorithms.
As demonstrated in Table 1, while all evaluated algorithms show improvement over the base SFT model, our RTO algorithm achieves superior performance across all benchmarks. Specifically, RTO outperforms PPO by achieving a 7.53% higher win rate in the AlpacaEval 2 LC benchmark and a 4.1% higher win rate in the Arena-Hard SC benchmark. These results highlight the effectiveness of incorporating token-wise reward (dense reward) into PPO training. Furthermore, when compared to the leading preference learning baselines, RTO demonstrates improvements of 2 and 4 points in the AlpacaEval 2 LC benchmark and the Arena-Hard SC benchmark, respectively.
Metric Method
SFT DPO R-DPO SimPO TDPO PPO RTO
AE (LC) 13.22 17.40 18.34 25.46 20.13 19.47 27.00
AE (WR) 8.58 12.23 12.03 20.20 11.97 12.89 22.45
AH (SC) 9.2 13.2 14.2 14.5 13.2 16.2 20.3
AH (WR) 8.9 13.8 14.1 15.2 12.3 15.6 21.4
Table 1: AlpacaEval 2 (AE) and Arena-Hard (AH) results.
5.2 In-depth Analysis of RTO Performance
In this subsection, we provide a detailed analysis of RTO. First, we examine the influence of reward granularity by comparing the performance of RL training based on different reward types. Next, we demonstrate that the token-wise reward (provided by DPO) primarily serves as reward shaping, which contributes significantly to RTO’s success. Finally, we investigate RTO’s sample efficiency to support our theoretical findings.
Method AlpacaEval 2 Arena-Hard
LC WR SC WR
RTO 27.00 22.45 20.3 21.4
Semi-RTO 23.77 19.17 19.0 19.7
DDPO 21.09 13.06 13.1 12.1
RS-PPO 27.52 21.69 19.2 19.9
Figure 3: Benchmark results of ablations studies.
The Influence of Reward Granularity.
We analyze three reward granularity settings by redistributing token-level rewards used in RTO: (i) RTO, where rewards are assigned to each token; (ii) Semi-RTO, where the rewards of all tokens in each sentence are reassigned to their delimiter, and (iii) DDPO, where all rewards are delayed and assigned to the EoS token. Table 3 and Figure 4(a) clearly demonstrates that denser rewards lead to better performance.
Reward Shaping via DPO Reward is the Key to RTO’s Success.
We demonstrate that the superior performance of RTO is not primarily due to replacing the reward model trained with MLE with the implicit reward from DPO. Instead, its advantage lies in its role as a reward-shaping mechanism. To illustrate this, we compare RTO and DPPO with another setup, RS-PPO, where the reward matches $r_{\mathrm{rto}}$ in (4.7), except for subtracting the DPO implicit reward $\beta_{1}\log\frac{\pi_{\mathrm{dpo}}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}$ from the last token. This adjustment results in a total reward equivalent to $r_{\mathrm{MLE}}(x,y)$, effectively employing the DPO token-wise implicit reward for reward shaping. From Table 3 and Figure 4(b), we observe that the main contribution of the DPO reward to improving RL training lies in reward shaping rather than altering the total reward through its exact value.
Sample Efficiency of RTO.
To validate the theoretical claims in Theorem 4.2, which guarantees the provable sample efficiency of RTO, and Proposition 3.2, which demonstrates that RTO is more efficient than PPO due to its use of token-wise rewards instead of sentence-wise rewards, we conducted experiments using only a fraction of the full dataset. These experiments evaluated the ability of RTO and PPO to learn an effective policy with limited data. As shown in Figure 4(c), RTO matches PPO’s performance using only about $1/8$ of the data and ultimately surpasses PPO’s final performance. Additionally, RTO exhibits superior data scaling behavior compared to PPO — RTO continues to improve with more data, while PPO’s performance saturates early.
((a)) Effect of different reward granularity.
((b)) Effect of reward shaping.
((c)) Data scaling behavior of PPO and RTO.
Figure 4: (a) and (b) show the obtained reward of $r_{\text{MLE}}$ throughout training. (c) shows the AlpacaEval 2 performance of PPO and RTO when trained on fractions of samples.
Additional Experiments on Summarization Task.
Due to space limit, we defer additional experiments on the summarization task (Völske et al., 2017) in Appendix E.